OpenStudy (anonymous):

Jackson wakes up at 8:00 a.m. and takes a walk for half an hour. He then spends 45 minutes bathing and eating breakfast. After watching the morning news on TV for a while, Jackson leaves for school at 9:30 a.m. How many minutes does Jackson spend watching the news?

OpenStudy (experimentx):

15 min

OpenStudy (anonymous):

45 mins

OpenStudy (king):

15 min

OpenStudy (anonymous):

Too many

OpenStudy (anonymous):

15 or 45 mins

OpenStudy (callisto):

In 1.5 hours you have (60+30 = 90) mins You've spent 45 mins bathing and eating breakfast and the remaining time on watching TV so time spent on watching TV = 90 -45 = 45mins